OpenStudy (anonymous):

Rewrite the following equation into slope-intercept form: 14x + 2y = 8 A. 2y = 14x + 8 B. 14x + 2y - 8 = 0 C. y = -7x + 4 D. x = 1/7y + 4/7 @iGreen. @Loser66

OpenStudy (anonymous):

Slope-intercept form is y = mx + b. First we subtract 14x to both sides: 14x + 2y = 8 2y = -14x + 8 Now we divide 2 to all terms. What's -14 / 2?

OpenStudy (anonymous):

-2 @iGreen.

OpenStudy (anonymous):

Help me plzz @shaadykhamlichi

OpenStudy (shaadykhamlichi):

ok

OpenStudy (anonymous):

No, it's not -2..check again.

OpenStudy (anonymous):

@tiffany__

OpenStudy (anonymous):

I Think its c

OpenStudy (anonymous):

Yep, C is correct. -14 / 2 = -7 8 / 2 = 4 So we have: y = -7x + 4

OpenStudy (anonymous):

@tiffany__